A car lift in a service station is used to lift a car of mass 2200 kilograms. If the radius of the output piston is 8 centimeters and the radius of the input piston is 3 centimeters, find the force needed to be supplied to the input piston in order to lift the car

1
Expert's answer
2021-12-13T16:55:08-0500

The force to the input piston is


"F=mg\frac{r_i^2}{r_o^2}=2200·9.8·\frac{3^2}{8^2}=3032\text{ N}."